### v2

# Self-Supervised Enhancement of Forward-Looking Sonar Images: Bridging Cross-Modal Degradation Gaps through Feature Space Transformation and Multi-Frame Fusion

Abstract
Enhancing forward-looking sonar images is critical for accurate underwater target detection. Current deep learning methods mainly rely on supervised training with simulated data, but the difficulty in obtaining high-quality real-world paired data limits their practical use and generalization. Although self-supervised approaches from remote sensing partially alleviate data shortages, they neglect the cross-modal degradation gap between sonar and remote sensing images. Directly transferring pretrained weights often leads to overly smooth sonar images, detail loss, and insufficient brightness. To address this, we propose a feature-space transformation that maps sonar images from the pixel domain to a robust feature domain, effectively bridging the degradation gap. Additionally, our self-supervised multi-frame fusion strategy leverages complementary inter-frame information to naturally remove speckle noise and enhance target-region brightness. Experiments on three self-collected real-world forward-looking sonar datasets show that our method significantly outperforms existing approaches, effectively suppressing noise, preserving detailed edges, and substantially improving brightness, demonstrating strong potential for underwater target detection applications.
Index Terms:
Forward-Looking Sonar, Self-Supervised Learning, Multi-Frame Fusion, Cross-Modal Degradation Gap.
I Introduction
Forward-looking sonar (FLS), with its exceptional underwater imaging capability, has become an indispensable tool in underwater target detection, environmental perception, and autonomous navigation tasks[1, 2, 3, 4, 5]. However, in complex underwater environments, sound waves are affected by factors such as water boundary reflection, multipath interference, and sound absorption by target materials, leading to significant degradation of sonar images. Common issues include speckle noise, uneven brightness, and brightness artifacts[6, 7, 8]. These challenges hinder tasks like target detection and 3D reconstruction, limiting the practical application and development of FLS technology in real-world underwater scenarios[9, 10, 11]. Efficiently enhancing forward-looking sonar images in such environments is a key issue in underwater robotics and marine exploration[12, 13, 14].
Figure 1: (a) Bridging the Cross-Modal Gap. The degradation gap between multispectral remote sensing (RS) and forward-looking sonar (FLS) images is effectively reduced by mapping both modalities into a unified feature space using wavelet scattering transform (WST). (b) Enhanced Results with Our Method. Compared to the overly smoothed baseline, our method preserves clear target contours and distinct foreground-background boundaries.
Sonar image enhancement methods can be divided into traditional and deep learning-based approaches. Traditional methods[15, 16, 17] improve image quality through filtering, contrast adjustment, and gray-level transformations but struggle to enhance severely degraded images due to limited information. Deep learning methods, especially supervised learning[18, 19, 20, 21], have made significant progress by learning the mapping from degraded to high-quality images using a large number of high-quality paired training samples. However, obtaining such paired sonar image data in underwater environments is challenging. As a result, existing research[22, 23] relies heavily on simulated data for training, leading to poor generalization to real-world data. Self-supervised learning, which does not require reference images, has emerged as an effective solution for enhancing low-quality sonar images in real scenarios. However, existing methods[24, 25] have not validated their enhancement effects on more complex targets and do not fully address the cross-modal degradation gap between multispectral remote sensing[26, 27, 28] and forward-looking sonar images, as shown in the upper part of Figure 1(a). While both are affected by speckle noise, their imaging mechanisms and noise characteristics differ significantly. Multispectral remote sensing images[29, 30] are approximated by pinhole imaging, suffer from low contrast and blurring due to uneven illumination and atmospheric attenuation, while forward-looking sonar images are distance-based[1, 2], with multipath effects from water scattering resulting in weak contrast and uneven brightness. Thus, directly transferring models trained on remote sensing images does not work effectively for sonar images, leading to issues such as overly smooth enhancement results, missing details, and insufficient brightness, as shown in the upper part of Figure 1(b).
To address these issues, we propose a self-supervised enhancement method for forward-looking sonar images, based on feature space transformation and multi-frame fusion. First, we use traditional feature extraction techniques (such as HOG, Canny, Haar, and Wavelet features)[31, 32, 33, 15, 34] to map the raw sonar image from a low-quality pixel space to a more robust feature space, reducing the cross-modal degradation gap between remote sensing and sonar images, as shown in the lower part of Figure 1(a). Second, a self-supervised multi-frame fusion framework is adopted to fully utilize complementary information between multiple frames for natural denoising and target enhancement. This eliminates the need for high-quality paired training data, improving the model’s generalization and practical applicability in underwater environments. Compared to previous self-supervised methods, our approach effectively removes noise, preserves target edge details, and significantly enhances brightness in target regions, as shown in the lower part of Figure 1(b).
The main contributions of this paper are summarized as follows:
  • Proposes an effective transfer learning approach for sonar image processing, mapping sonar images to a robust feature space to reduce the cross-modal degradation gap with other domains (e.g., remote sensing images).
  • Designs a self-supervised enhancement framework for forward-looking sonar images based on multi-frame fusion, achieving significant enhancement effects without requiring reference images.
  • Constructs a multi-view forward-looking sonar image dataset with three types of targets made from different materials, exhibiting diverse degradation characteristics, and providing accurate pose parameters for each image. Experiments on this dataset show that our method outperforms existing mainstream methods in terms of visual quality and quantitative metrics.
Figure 2: Overview of the proposed method. The main idea is that the model integrates a feature space transformation module and a multi-frame image fusion network to bridge the cross-modal degradation gap and enhance the quality of forward-looking sonar images through self-supervised learning.The original images, heavily degraded by speckle noise and low target-background contrast, are significantly improved after enhancement, exhibiting greater brightness, sharper target contours, and well-preserved edge details.
II Related Work
II-A Quality Degradation of Forward-Looking Sonar Images and Traditional Enhancement Methods
Forward-looking sonar images often experience quality degradation in complex underwater environments[35, 36, 37, 38], characterized by speckle noise, uneven illumination, and brightness artifacts due to multipath effects and acoustic absorption by target materials[39, 40]. Among these, speckle noise is multiplicative interference from multiple acoustic paths[6, 7], greatly reducing target-background contrast. Traditional enhancement methods primarily aim at suppressing speckle noise[8, 9].
These traditional methods are generally categorized into spatial-domain and transform-domain approaches. Spatial-domain methods, such as Gaussian smoothing and median filtering, reduce speckle noise but often lead to over-smoothing, causing blurred details and decreased brightness[17]. Transform-domain methods, including wavelet and Fourier transforms[15, 16], suppress noise and enhance details within the transformed space. However, while effectively removing high-frequency noise, these approaches frequently reduce overall image contrast and brightness.
In addition, these methods incorporate various handcrafted feature extraction techniques, such as HOG, Canny edge detection, Haar-like features, and wavelet scattering transform[31, 32, 33, 15, 34]. By extracting key structural information, these techniques transform images from the pixel domain to a more robust feature space, improving image quality under noisy conditions. We argue that such feature-space transformations effectively reduce cross-modal degradation gaps and provide a practical framework for transferring enhancement models trained in other domains to sonar image processing tasks.
II-B Deep-Learning-Based Sonar Image Enhancement Methods
With the rapid advancement of deep learning, methods for enhancing and denoising forward-looking sonar images have continually evolved. Central to these methods is the challenge of generating high-quality sonar images for supervised training. For example, compressive sensing (CS) techniques and autoencoders have been applied for feature extraction in denoising, with Gaussian noise added to the original sonar images as demonstrated by [18, 19]. However, both methods overlook that actual sonar image noise is multiplicative, not additive. UFIDNet[20] simulates this multiplicative noise by multiplying clear images with gamma-distributed noise, using logarithmic transformation to separate intensity and reflection, converting speckle noise into additive noise that is easier to remove. Furthermore, [21]converts optical RGBD images into 3D point clouds and uses Blender to simulate the sonar imaging process in 3D space, generating images with background noise, propagation noise, multipath effects, and azimuthal artifacts. While these methods have made progress, they rely on clear real or simulated data for supervised training, and acquiring high-quality paired data in real underwater scenarios remains challenging, limiting their generalization in practical applications.
To reduce reliance on paired data, methods combining traditional transform-domain techniques and deep learning have emerged. For instance, FDBANet [41] introduces a frequency-domain denoising attention module using FFT to convert sonar images to the frequency domain, filtering noise and enhancing high-frequency features. While this method works well for low-resolution images, its enhancement is limited due to the lack of a data-driven learning mechanism. Self-supervised learning[42, 43, 44, 45], however, does not require reference images and directly utilizes real low-resolution sonar images for training. The SEGSID framework[24] enhances local receptive fields and extracts global semantic information, overcoming traditional blind spot network (BSN) limitations in handling speckle noise with strong spatial correlation. Yet, it uses only single-image information, limiting its ability to enhance image resolution. In contrast, multi-frame fusion methods leverage complementary information across frames, naturally removing noise and improving resolution[25]. This method boosts the brightness of small target areas by fine-tuning a pre-trained remote sensing model but does not fully address the cross-modal degradation gap between multispectral remote sensing and sonar images[34]. In regions with complex structures or severe degradation, the enhanced image can become overly smooth, losing edge details and affecting subsequent target detection accuracy.
III Method
III-A Framework Overview
As illustrated in Figure 2, the self-supervised forward-looking sonar image enhancement method proposed in this study consists of two main components: the feature space transformation module and the multi-frame image fusion enhancement network.
Both multispectral remote sensing and forward-looking sonar images suffer from speckle noise and low resolution, but a cross-modal degradation gap exists between them[46, 47]. Directly transferring pre-trained models from multispectral data to sonar images often leads to overly smooth, detail-lacking, and low-brightness results. To address this, we designed the feature space transformation module. By applying handcrafted filtering features, such as wavelet scattering transforms, we convert the original sonar images from a low-quality pixel domain into a more robust feature space. This transformation effectively reduces the degradation gap between remote sensing and sonar images, enabling better adaptation of pre-trained models to real sonar data.
The extracted features from both the feature and pixel spaces are concatenated and passed through convolution operations before being fed into the multi-frame image fusion enhancement network. Using self-supervised learning, complementary information from consecutive frames is integrated to enhance image details and brightness. As the sonar data includes accurate pose information, no additional image registration is needed before multi-frame fusion. This framework minimizes dependence on paired supervised data, achieving significantly improved enhancement results on three self-collected real-world forward-looking sonar dataset compared to existing methods.
III-B Feature Space Transformation Module
To bridge the cross-modal degradation gap between multispectral remote sensing and forward-looking sonar images, we designed the feature space transformation module. This module maps the original sonar images from pixel space to a more robust feature space using handcrafted filtering features, which enhances noise suppression and captures structural information, thus narrowing the gap and providing consistent input for the subsequent self-supervised multi-frame fusion enhancement.
Figure 3: Visualization of handcrafted features. WST is selected for its balanced performance in denoising and detail retention (see Ablation Study).
The module applies a series of predefined filters to the original single-channel sonar images $I$, extracting robust feature space information. Figure 3 shows the visual results of these features:
$G_{i}(I)=\Phi_{i}(I),i\in\{\text{HOG},\text{Canny},\text{GRE},\text{Haar},\text{WST}\}$ (1)
where ${\Phi_{i}}$ denotes their respective filtering transformation functions, and ${G_{i}(I)}$ denotes the extracted filtering features.Next, we concatenate the original image $I$ with the extracted features ${G_{i}(I)}$, creating an enhanced input representation:
$I_{aug}=\text{concat}(I,\{G_{i}(I)\})$ (2)
To integrate these multi-channel features, we use $1\times 1$ convolution operations ${\rm{\Psi}}(\cdot)$ to compress and unify the channels, producing the final fused output:
$I_{fused}=\Psi(I_{aug})$ (3)
As shown in Figure 1, the original forward-looking sonar image (red) and multispectral remote sensing image (green) exhibit a cross-modal degradation gap. The feature space transformation module serves as a bridge, converting the data from the heterogeneous pixel space to a unified feature space (yellow), significantly reducing the gap and enabling better adaptation of pre-trained models from remote sensing data to our self-supervised multi-frame fusion network.
III-C Multi-Frame Image Fusion Network
Figure 4: Illustration of the multi-frame fusion network. The proposed network recursively fuses aligned forward-looking sonar frames in a pairwise manner—each step combines the current state with a new frame—until all K frames are aggregated. A median reference image guides self-supervised enhancement through downsampling and gradient consistency losses, enabling effective speckle noise suppression and high-resolution output (from 376×376 to 800×800).
Forward-looking sonar systems typically capture images at high frame rates, which allows for the effective use of complementary information between consecutive frames to naturally eliminate speckle noise. As shown in Figure 4, our designed multi-frame image fusion network does not require an additional registration module because the collected sonar images come with accurate pose information.
The overall architecture comprises three modules: encoding, fusion, and decoding. Let the input low-resolution image sequence be denoted by
$J=\{J(1),J(2),\ldots,J(K)\}$ (4)
where ${J^{(i)}}$ represents the $i^{\text{th}}$ frame. The encoding module ${E(\cdot)}$ extracts features from each frame, mapping the sequence to a high-dimensional feature space; the fusion module ${F(\cdot)}$ integrates these features into a global encoding ${Z=F(E(J))}$ and the decoding module ${D(\cdot)}$ upsamples and reconstructs the global encoding ${Z}$ to generate the final super-resolution image ${Y}$, namely
$Y=D(F(E(J)))$ (5)
Since it is challenging to acquire high-quality sonar images in real underwater scenarios, we follow [25] and adopt a downsampling mean squared error (MSE) loss (denoted as ${{\cal L}_{DS}}$) for self-supervised training. This loss ensures that the enhanced super-resolution image remains globally consistent, on a pixel level, with the original low-resolution image. Let the downsampling function be ${H(\cdot)}$, and define ${J^{*}}$ as the median image, which must satisfy:
$H(Y)\approx J^{*}$ (6)
$J^{*}(u,v)=\operatorname{median}\{J^{(1)}(u,v),J^{(2)}(u,v),\ldots,J^{(M)}(u,v)\}$ (7)
where ${(u,v)}$ denotes the pixel coordinates in the image. Optimization is performed using the downsampling MSE loss:
$L_{down}=\|H(Y)-J^{*}\|_{2}^{2}$ (8)
However, because the downsampling loss primarily focuses on global pixel information, it tends to overly smooth high-frequency details, thereby weakening key features in target regions and affecting subsequent target detection. To address this issue, we introduce a local consistency loss and a gradient consistency loss to ensure that the enhanced super-resolution image preserves the local variation trends of the low-resolution image. By comparing the differences in adjacent pixel regions, we can effectively retain high-frequency details and avoid over-smoothing:
$L_{\mathrm{con}}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\left(\left|Y_{p}-Y_{q}\right|-\left|J_{p}^{*}-J_{q}^{*}\right|\right)$ (9)
$L_{\mathrm{grad}}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\left(\left|\nabla(Y_{p})-\nabla(Y_{q})\right|-\left|\nabla(J_{p}^{*})-\nabla(J_{q}^{*})\right|\right)$ (10)
[…]
The final loss function ${L_{\mathrm{total}}}$ is expressed as:
$L_{\rm total}=L_{\rm down}+\lambda_{\rm con}L_{\rm con}+\lambda_{\rm grad}L_{\rm grad}$ (11)
where ${\lambda_{\mathrm{con}}}$ and ${\lambda_{\mathrm{grad}}}$ are the hyperparameters that balance each component of the loss function.
IV Experiments
In this section, we demonstrate that our method outperforms existing approaches [15, 25] on a real-world forward-looking sonar dataset containing three target types. Our experiments, conducted on a 24GB NVIDIA GeForce GTX 4090, show that the accurate pose information in the sonar images eliminates the need for additional registration before multi-frame fusion, allowing 16 images to be fused in just 0.3 seconds during inference.
IV-A Forward-Looking Sonar Datasets
Figure 5: Schematic diagram of the forward-looking sonar acquisition system and representative target images.(a) Custom-built water tank for controlled data collection; (b) C900-II sonar capable of capturing high-frame-rate sequences; (c) 360° rotating guide rail enables accurate pose estimation and multi-view imaging without additional registration; (d) three designed targets with distinct materials—rubber tire, metal torpedo model, and GRP conical frustum.
We built a standard data collection pool for forward-looking sonar (see Figure 5) using a C900-II unit operating at 2.7 MHz, with a horizontal beamwidth of 50° and a vertical beamwidth of 15°. The sonar is fixed 1.5 m from the guide rail cantilever and tilted 15° downward relative to the rotating platform. During acquisition, the system performs a 360° rotation around the target, with the guide rail rotation angle recorded continuously. Combining this rotation data with the initial sonar pose enables calculation of the sensor’s pose at each instance, eliminating the need for additional image registration prior to multi-frame fusion.
The dataset comprises multi-frame forward-looking sonar images capturing three target types from various materials and viewpoints: rubber tires (exhibiting severe acoustic absorption), metal torpedo models, and glass-reinforced plastic (GRP) conical frustum. Each target class contains 3,200 images grouped into sets of 16 frames. We randomly selected 180 groups for training and 10 groups each for validation and testing. All images have been converted from polar to Cartesian coordinates at a resolution of 376 × 376.
IV-B Evaluation Metrics for Enhancement Performance
In underwater environments, high-quality reference images are scarce, making common metrics for optical image enhancement, such as PSNR, unsuitable. We adopt methods from [24, 25] to evaluate the denoising performance and the preservation of high-frequency details in sonar image enhancement. Standard deviation (STD) is used to assess image smoothness—lower STD indicates stronger noise suppression—while average gradient (AG) measures high-frequency detail, with higher AG reflecting richer detail.
$STD=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(I_{i}-\mu\right)^{2}}$ (12)
$AG=\frac{1}{N}\sum_{i=1}^{N}\sqrt{\left(I_{x,i}-I_{x,i-1}\right)^{2}+\left(I_{y,i}-I_{y,i-1}\right)^{2}}$ (13)
where ${I_{i}}$ denotes the grayscale value of the $i^{\text{th}}$ pixel, ${I_{x,i}}$ and ${I_{y,i}}$ represent the horizontal and vertical gradients, respectively, ${\mu}$ is the mean value of the image, and ${N}$ is the total number of pixels.
IV-C Quantitative and Visual Evaluation of Enhancement Results
Table I presents a quantitative comparison across three representative target types: torpedo model, tire, and conical frustum. Our method consistently achieves the best results in both standard deviation (STD) and average gradient (AG), demonstrating strong capabilities in suppressing noise and preserving structural details.
For the torpedo model, the STD is reduced from 17.02 to 13.14, surpassing wavelet-based (21.59) and HighRes-based (14.58) methods. AG improves from 4.93 to 5.92, indicating sharper edges and clearer target contours.
In the tire category—challenging due to strong acoustic absorption—our method delivers the most substantial improvement, reducing STD from 58.79 to 37.68 and boosting AG from 8.09 to 14.71. This highlights its robustness in handling severe degradation while maintaining fine details.
For the conical frustum, our method again outperforms all baselines, achieving the lowest STD (23.24) and highest AG (13.66), confirming its adaptability to varying materials and geometries.
Overall, the proposed method demonstrates consistently superior performance across all scenarios. By combining feature space transformation with multi-frame fusion, it effectively balances speckle noise suppression and detail preservation—crucial for accurate underwater target detection in forward-looking sonar imagery.
These quantitative improvements are further supported by visual comparisons. Figure 6 presents enhancement results for the same three targets, with each row showing enlarged views of the target regions for detailed inspection. Existing methods either amplify background noise or blur target contours, leading to poor contrast and reduced structural clarity. In contrast, our method enhances target intensity and edge sharpness while maintaining a clean background, closely aligning with the improvements observed in Table I.
Furthermore, Figure 8 illustrates enhancement results under different sonar viewpoints. The original degraded images exhibit strong speckle noise and weak contrast. After applying feature space transformation and multi-frame fusion, our method produces significantly brighter target regions and clearer boundaries, demonstrating robustness to viewpoint variation and structural deformation.
In summary, these quantitative and visual results together validate the robustness and effectiveness of our proposed method. By integrating feature space transformation with multi-frame fusion, the approach consistently achieves superior performance across diverse targets and degradation conditions. It excels in suppressing speckle noise, preserving fine structural details, and maintaining stable enhancement across varying viewpoints—highlighting its strong generalizability and practical value for real-world forward-looking sonar image enhancement.
TABLE I: Enhancement performance comparison for three different target types.
(a) Torpedo model
Fusion Method Original Wavelet [15] Highres [25] Proposed
STD $\downarrow$ 17.02 21.59 14.58 13.14
AG $\uparrow$ 4.93 5.06 3.70 5.92
(b) Tire
Fusion Method Original Wavelet [15] Highres [25] Proposed
STD $\downarrow$ 58.79 45.96 48.80 37.68
AG $\uparrow$ 8.09 9.61 10.62 14.71
(c) Conical frustum
Fusion Method Original Wavelet [15] Highres [25] Proposed
STD $\downarrow$ 30.41 30.42 28.83 23.24
AG $\uparrow$ 9.99 7.99 10.32 13.66
Figure 6: Visual comparison of enhancement performance across three target categories. Each row shows enlarged target regions (torpedo model, tire, and conical frustum) for detailed comparison. Existing methods either amplify noise or over-smooth targets, leading to blurred edges and poor contrast. In contrast, our method enhances brightness, preserves clear contours, and suppresses background noise effectively.
TABLE II: Enhancement performance of different feature operators for the torpedo model target.
Input STD $\downarrow$ AG $\uparrow$
FLR 17.016 4.932
FLR + HOG 27.338 4.809
FLR + Canny 8.978 1.552
FLR + GRE 16.277 3.131
FLR + HAAR 10.348 1.683
FLR + WST 13.143 5.918
Figure 7: Comparison of enhancement results using Canny and WST transformations. WST preserves high-frequency details with stronger target contrast and clearer boundaries, while Canny produces oversmoothed results with blurred edges, limiting detection accuracy.
IV-D Ablation Study
To further assess the contribution of the proposed feature space transformation module, we conducted ablation studies within our multi-frame self-supervised enhancement framework, incorporating several widely used feature operators in sonar image processing. These include Histogram of Oriented Gradients (HOG), Canny edge detection, Gray-Level Co-occurrence Matrix (GRE), Haar wavelet features (HAAR), and Wavelet Scattering Transform (WST).
Figure 8: Enhancement results across varying sonar viewpoints. (a) Original images exhibit heavy speckle noise and low contrast across perspectives. (b) Feature space transformation reduces cross-modal discrepancies and reinforces structural consistency. (c) The proposed method enhances brightness and edge clarity, achieving robust performance under viewpoint changes.
Table II reports the quantitative results using standard deviation (STD) and average gradient (AG). The results highlight notable performance differences. Canny achieved the lowest STD, indicating effective noise suppression. However, its low AG suggests over-smoothing and the loss of critical high-frequency edge details.
In contrast, WST delivered the most balanced performance, combining strong noise reduction with superior detail preservation. This is largely due to its multi-scale analysis capability, which enables robust extraction and retention of structural features, particularly in low-quality forward-looking sonar images.
Figure 7 further visualizes the enhancement results of Canny and WST. The WST-based output exhibits higher overall intensity, clearer boundaries, and richer detail, while the Canny-based result shows evident blurring and loss of edge information. Such high-frequency detail and contrast enhancement are essential for accurate underwater target detection.
In conclusion, the ablation study demonstrates that WST provides the optimal trade-off between denoising and detail retention, making it the most effective operator for bridging the cross-modal degradation gap. As such, WST is selected as the default feature transformation module in our final approach.
V Conclusion
This paper presents a self-supervised enhancement method for forward-looking sonar images, combining feature space transformation with multi-frame fusion. The proposed transformation module effectively bridges the cross-modal degradation gap between sonar and other domains (e.g., multispectral remote sensing), enabling a practical transfer learning strategy. A self-supervised fusion network leverages the high frame rate of sonar systems to integrate complementary frame information, naturally suppressing noise while preserving high-frequency details. Experiments on three real-world, heavily degraded sonar datasets show that the method outperforms existing approaches in both quantitative metrics and visual quality. With an inference time of 0.3 seconds for 16 frames, the method is also efficient and well-suited for deployment on underwater vehicles to enhance target detection.
VI Acknowledgment
This work was supported by the National Natural Science Foundation of China (No. 62401601).

### v3

@@
-Self-Supervised Enhancement of Forward-Looking Sonar Images: Bridging Cross-Modal Degradation Gaps through Feature Space Transformation and Multi-Frame Fusion
+Reference-Free Enhancement of Forward-Looking Sonar Images: Bridging Cross-Modal Degradation Gaps Through Deformable Wavelet Scattering Transform and Multi-Frame Fusion
 Abstract
-Enhancing forward-looking sonar images is critical for accurate underwater target detection.
-Current deep learning methods mainly rely on supervised training with simulated data, but the difficulty in obtaining high-quality real-world paired data limits their practical use and generalization.
-Although self-supervised approaches from remote sensing partially alleviate data shortages, they neglect the cross-modal degradation gap between sonar and remote sensing images.
-Directly transferring pretrained weights often leads to overly smooth sonar images, detail loss, and insufficient brightness.
-To address this, we propose a feature-space transformation that maps sonar images from the pixel domain to a robust feature domain, effectively bridging the degradation gap.
-Additionally, our self-supervised multi-frame fusion strategy leverages complementary inter-frame information to naturally remove speckle noise and enhance target-region brightness.
-Experiments on three self-collected real-world forward-looking sonar datasets show that our method significantly outperforms existing approaches, effectively suppressing noise, preserving detailed edges, and substantially improving brightness, demonstrating strong potential for underwater target detection applications.
+Enhancing forward-looking sonar images is crucial for effective underwater target detection.
+Existing deep learning approaches typically depend on supervised training using simulated data, but their performance and generalization are limited by the scarcity of real-world paired datasets.
+Although recent reference-free methods have mitigated reliance on paired data, they overlook the inherent cross-modal degradation gap between sonar and remote sensing images.
+Direct application of pretrained remote-sensing models to sonar data frequently yields overly smoothed images with lost details and inadequate brightness.
+To overcome this challenge, we propose a novel Deformable Wavelet Scattering Transform (WST) Feature Bridge, which adaptively projects sonar images into a multi-scale, translation-invariant feature domain, effectively narrowing the cross-modal degradation gap.
+Additionally, we develop a reference-free multi-frame fusion strategy that exploits complementary inter-frame information to suppress speckle noise and enhance target visibility without external ground-truth.
+Experiments conducted on three sonar datasets demonstrate that the proposed framework substantially surpasses existing reference-free methods in terms of noise suppression, edge preservation, and brightness enhancement, highlighting its practical effectiveness for underwater detection scenarios.
 Index Terms:
-Forward-Looking Sonar, Self-Supervised Learning, Multi-Frame Fusion, Cross-Modal Degradation Gap.
+Forward-Looking Sonar, reference-free, Multi-Frame Fusion, Cross-Modal Degradation Gap.
 I Introduction
-Forward-looking sonar (FLS), with its exceptional underwater imaging capability, has become an indispensable tool in underwater target detection, environmental perception, and autonomous navigation tasks[1, 2, 3, 4, 5].
-However, in complex underwater environments, sound waves are affected by factors such as water boundary reflection, multipath interference, and sound absorption by target materials, leading to significant degradation of sonar images.
-Common issues include speckle noise, uneven brightness, and brightness artifacts[6, 7, 8].
-These challenges hinder tasks like target detection and 3D reconstruction, limiting the practical application and development of FLS technology in real-world underwater scenarios[9, 10, 11].
-Efficiently enhancing forward-looking sonar images in such environments is a key issue in underwater robotics and marine exploration[12, 13, 14].
+Forward-looking sonar (FLS) provides unique underwater imaging capabilities, becoming essential in underwater target detection, environmental perception, and autonomous navigation tasks [1, 2, 3, 4, 5].
+However, complex underwater environments significantly degrade sonar images due to multipath scattering, acoustic absorption by target materials, and reflections from environmental boundaries.
+Such effects typically introduce severe speckle noise, brightness inconsistency, and various imaging artifacts [6, 7, 8], hindering accurate target detection and reliable 3D reconstruction, thereby limiting practical deployment of FLS in real-world scenarios [9, 10, 11].
 Figure 1: (a) Bridging the Cross-Modal Gap.
-The degradation gap between multispectral remote sensing (RS) and forward-looking sonar (FLS) images is effectively reduced by mapping both modalities into a unified feature space using wavelet scattering transform (WST).
-(b) Enhanced Results with Our Method.
+The degradation gap between multispectral remote sensing (RS) and forward-looking sonar (FLS) images is effectively narrowed by mapping both modalities into a unified feature space via our Deformable WST Feature Bridge; (b) Enhanced Results with Our Method.
 Compared to the overly smoothed baseline, our method preserves clear target contours and distinct foreground-background boundaries.
-Sonar image enhancement methods can be divided into traditional and deep learning-based approaches.
-Traditional methods[15, 16, 17] improve image quality through filtering, contrast adjustment, and gray-level transformations but struggle to enhance severely degraded images due to limited information.
-Deep learning methods, especially supervised learning[18, 19, 20, 21], have made significant progress by learning the mapping from degraded to high-quality images using a large number of high-quality paired training samples.
-However, obtaining such paired sonar image data in underwater environments is challenging.
-As a result, existing research[22, 23] relies heavily on simulated data for training, leading to poor generalization to real-world data.
-Self-supervised learning, which does not require reference images, has emerged as an effective solution for enhancing low-quality sonar images in real scenarios.
-However, existing methods[24, 25] have not validated their enhancement effects on more complex targets and do not fully address the cross-modal degradation gap between multispectral remote sensing[26, 27, 28] and forward-looking sonar images, as shown in the upper part of Figure 1(a).
-While both are affected by speckle noise, their imaging mechanisms and noise characteristics differ significantly.
-Multispectral remote sensing images[29, 30] are approximated by pinhole imaging, suffer from low contrast and blurring due to uneven illumination and atmospheric attenuation, while forward-looking sonar images are distance-based[1, 2], with multipath effects from water scattering resulting in weak contrast and uneven brightness.
-Thus, directly transferring models trained on remote sensing images does not work effectively for sonar images, leading to issues such as overly smooth enhancement results, missing details, and insufficient brightness, as shown in the upper part of Figure 1(b).
-To address these issues, we propose a self-supervised enhancement method for forward-looking sonar images, based on feature space transformation and multi-frame fusion.
-First, we use traditional feature extraction techniques (such as HOG, Canny, Haar, and Wavelet features)[31, 32, 33, 15, 34] to map the raw sonar image from a low-quality pixel space to a more robust feature space, reducing the cross-modal degradation gap between remote sensing and sonar images, as shown in the lower part of Figure 1(a).
-Second, a self-supervised multi-frame fusion framework is adopted to fully utilize complementary information between multiple frames for natural denoising and target enhancement.
-This eliminates the need for high-quality paired training data, improving the model’s generalization and practical applicability in underwater environments.
-Compared to previous self-supervised methods, our approach effectively removes noise, preserves target edge details, and significantly enhances brightness in target regions, as shown in the lower part of Figure 1(b).
-The main contributions of this paper are summarized as follows:
-  • Proposes an effective transfer learning approach for sonar image processing, mapping sonar images to a robust feature space to reduce the cross-modal degradation gap with other domains (e.g., remote sensing images).
-  • Designs a self-supervised enhancement framework for forward-looking sonar images based on multi-frame fusion, achieving significant enhancement effects without requiring reference images.
-  • Constructs a multi-view forward-looking sonar image dataset with three types of targets made from different materials, exhibiting diverse degradation characteristics, and providing accurate pose parameters for each image.
-Experiments on this dataset show that our method outperforms existing mainstream methods in terms of visual quality and quantitative metrics.
+Current sonar image enhancement methods generally fall into two categories [12, 13, 14]: traditional filtering-based techniques and supervised deep-learning approaches.
+Traditional methods, including filtering, contrast adjustment, and grayscale transformations [15, 16, 17], often struggle to restore severely degraded sonar images effectively.
+Supervised deep-learning approaches [18, 19, 20, 21], while capable of achieving better enhancement, heavily depend on large-scale, high-quality paired datasets.
+Collecting such datasets in realistic underwater environments is notoriously challenging, which severely restricts their generalizability [22, 23].
+Recent reference-free enhancement methods [24, 25] circumvent this dependence but have yet to address the intrinsic cross-modal degradation differences between forward-looking sonar and multispectral remote sensing images [26, 27, 28].
+Although both modalities exhibit speckle noise, remote sensing images suffer from blur and low contrast due to uneven illumination and atmospheric attenuation under a pinhole camera model [29, 30], whereas sonar imaging is fundamentally distance-based, resulting in multipath scattering-induced brightness variations and weaker contrast [1, 2].
+Consequently, directly applying pretrained remote-sensing models to sonar enhancement frequently produces overly smoothed outputs with loss of essential details (Fig.
+1(a), (b)).
+To address these limitations, we propose a novel reference-free sonar enhancement pipeline that integrates a Deformable Wavelet Scattering Transform (WST) Feature Bridge and a multi-frame fusion network into an end-to-end framework.
+Specifically, our Deformable WST Feature Bridge employs learnable perturbations on wavelet scales and orientations—carefully selected via ablation among multiple handcrafted features [31, 32, 33, 15, 34]—to adaptively map sonar images into a robust, translation-invariant feature domain, significantly narrowing the cross-modal degradation gap.
+Subsequently, our multi-frame fusion network leverages complementary information from consecutive sonar frames, naturally suppressing speckle noise and enhancing target-region brightness without relying on external high-quality ground truth.
+Unlike previous methods, our unified approach simultaneously optimizes adaptive feature extraction and multi-frame fusion, resulting in sharper contours, richer details, and enhanced brightness in sonar imagery.
+The main contributions of this paper can be summarized as follows:
+  • We propose a Deformable WST Feature Bridge that adaptively aligns sonar image features with pretrained remote-sensing distributions by embedding learnable offsets into wavelet scales and orientations, effectively reducing the cross-modal degradation gap.
+  • We develop a reference-free multi-frame fusion framework for sonar image enhancement, utilizing inter-frame complementary information to suppress speckle noise and enhance target contrast without external reference data.
+  • We construct a carefully annotated sonar image dataset featuring three material-diverse targets (rubber tire, metal torpedo, GRP frustum) captured along circular and linear trajectories, exhibiting distinct degradation patterns (multipath scattering, acoustic absorption).
+Precise pose annotations eliminate the need for image registration and enable rigorous benchmarking.
+Experiments demonstrate superior performance of our method over state-of-the-art baselines both visually and quantitatively.
+Code and Dataset Release. The source code and the collected sonar dataset will be publicly released upon publication.
 Figure 2: Overview of the proposed method.
-The main idea is that the model integrates a feature space transformation module and a multi-frame image fusion network to bridge the cross-modal degradation gap and enhance the quality of forward-looking sonar images through self-supervised learning.The original images, heavily degraded by speckle noise and low target-background contrast, are significantly improved after enhancement, exhibiting greater brightness, sharper target contours, and well-preserved edge details.
+The model combines a Deformable WST Feature Bridge—with learnable wavelet scale and orientation offsets—and a multi-frame fusion network into one end-to-end pipeline, replacing the fixed WST extractor plus simple concat-and-$1\times 1$ convolution.
+This unified design closes the cross-modal degradation gap, reduces speckle noise, and boosts target brightness, producing images with crisper contours and richer edge detail.
 II Related Work
@@
 However, while effectively removing high-frequency noise, these approaches frequently reduce overall image contrast and brightness.
-In addition, these methods incorporate various handcrafted feature extraction techniques, such as HOG, Canny edge detection, Haar-like features, and wavelet scattering transform[31, 32, 33, 15, 34].
+In addition, these methods incorporate various handcrafted feature extraction techniques, such as Histogram of Oriented Gradients (HOG), Canny edge detection (Canny), Haar-like features (HAAR), Gaussian Radial Edges (GRE) and Wavelet Scattering Transform (WST) [31, 32, 33, 15, 34].
 By extracting key structural information, these techniques transform images from the pixel domain to a more robust feature space, improving image quality under noisy conditions.
@@
 To reduce reliance on paired data, methods combining traditional transform-domain techniques and deep learning have emerged.
-For instance, FDBANet [41] introduces a frequency-domain denoising attention module using FFT to convert sonar images to the frequency domain, filtering noise and enhancing high-frequency features.
+For instance, FDBANet [41] introduces a frequency-domain denoising attention module using Fast Fourier Transform (FFT) to convert sonar images to the frequency domain, filtering noise and enhancing high-frequency features.
 While this method works well for low-resolution images, its enhancement is limited due to the lack of a data-driven learning mechanism.
-Self-supervised learning[42, 43, 44, 45], however, does not require reference images and directly utilizes real low-resolution sonar images for training.
+reference-free methods [42, 43, 44, 45], however, does not require reference images and directly utilizes real low-resolution sonar images for training.
 The SEGSID framework [24] enhances local receptive fields and extracts global semantic information, overcoming traditional blind spot network (BSN) limitations in handling speckle noise with strong spatial correlation.
@@
 III-A Framework Overview
-As illustrated in Figure 2, the self-supervised forward-looking sonar image enhancement method proposed in this study consists of two main components: the feature space transformation module and the multi-frame image fusion enhancement network.
-Both multispectral remote sensing and forward-looking sonar images suffer from speckle noise and low resolution, but a cross-modal degradation gap exists between them[46, 47].
-Directly transferring pre-trained models from multispectral data to sonar images often leads to overly smooth, detail-lacking, and low-brightness results.
-To address this, we designed the feature space transformation module.
-By applying handcrafted filtering features, such as wavelet scattering transforms, we convert the original sonar images from a low-quality pixel domain into a more robust feature space.
-This transformation effectively reduces the degradation gap between remote sensing and sonar images, enabling better adaptation of pre-trained models to real sonar data.
-The extracted features from both the feature and pixel spaces are concatenated and passed through convolution operations before being fed into the multi-frame image fusion enhancement network.
-Using self-supervised learning, complementary information from consecutive frames is integrated to enhance image details and brightness.
-As the sonar data includes accurate pose information, no additional image registration is needed before multi-frame fusion.
-This framework minimizes dependence on paired supervised data, achieving significantly improved enhancement results on three self-collected real-world forward-looking sonar dataset compared to existing methods.
-III-B Feature Space Transformation Module
-To bridge the cross-modal degradation gap between multispectral remote sensing and forward-looking sonar images, we designed the feature space transformation module.
-This module maps the original sonar images from pixel space to a more robust feature space using handcrafted filtering features, which enhances noise suppression and captures structural information, thus narrowing the gap and providing consistent input for the subsequent self-supervised multi-frame fusion enhancement.
-Figure 3: Visualization of handcrafted features. WST is selected for its balanced performance in denoising and detail retention (see Ablation Study).
-The module applies a series of predefined filters to the original single-channel sonar images $I$, extracting robust feature space information. Figure 3 shows the visual results of these features:
-$G_{i}(I)=\Phi_{i}(I),i\in\{\text{HOG},\text{Canny},\text{GRE},\text{Haar},\text{WST}\}$ (1)
-where ${\Phi_{i}}$ denotes their respective filtering transformation functions, and ${G_{i}(I)}$ denotes the extracted filtering features.Next, we concatenate the original image $I$ with the extracted features ${G_{i}(I)}$, creating an enhanced input representation:
-$I_{aug}=\text{concat}(I,\{G_{i}(I)\})$ (2)
-To integrate these multi-channel features, we use $1\times 1$ convolution operations ${\rm{\Psi}}(\cdot)$ to compress and unify the channels, producing the final fused output:
-$I_{fused}=\Psi(I_{aug})$ (3)
-As shown in Figure 1, the original forward-looking sonar image (red) and multispectral remote sensing image (green) exhibit a cross-modal degradation gap.
-The feature space transformation module serves as a bridge, converting the data from the heterogeneous pixel space to a unified feature space (yellow), significantly reducing the gap and enabling better adaptation of pre-trained models from remote sensing data to our self-supervised multi-frame fusion network.
+As shown in Figure 2, our reference-free forward-looking sonar enhancement pipeline consists of two key modules: the Deformable WST Feature Bridge and the Multi-Frame Fusion Network.
+The Deformable WST Feature Bridge applies learnable offsets to wavelet scales and orientations, adaptively transforming each raw sonar frame into a robust, multi-scale feature representation that aligns with pretrained remote-sensing features.
+These per-frame feature tensors are then stacked into a sequence and passed to the Multi-Frame Fusion Network, which comprises an encoder, a fusion block, and a decoder.
+By integrating complementary information across consecutive frames, the fusion network suppresses speckle noise and reconstructs a high-resolution output.
+All components are learned end-to-end under a reference-free objective—combining downsampling consistency with local and gradient consistency losses—and leverage accurate pose metadata to eliminate any need for explicit registration.
+III-B Deformable WST Feature Bridge
+To effectively bridge the cross-modal degradation gap between multispectral remote sensing and forward-looking sonar imagery, we introduce the Deformable Wavelet Scattering Transform (WST) Feature Bridge.
+This module adaptively transforms raw sonar images into a robust, unified feature representation tailored specifically for sonar image enhancement.
+Let $I\in\mathbb{R}^{H\times W}$ denote the input sonar image. We first obtain a set of handcrafted feature maps:
+$G_{i}(I)=\Phi_{i}(I),\quad i\in\{\mathrm{HOG},\mathrm{Canny},\mathrm{GRE},\mathrm{Haar},\mathrm{WST}\},$ (1)
+where each $\Phi_{i}(\cdot)$ applies the corresponding filter and $G_{i}(I)\in\mathbb{R}^{H\times W}$.
+Based on our ablation study, the Wavelet Scattering Transform (WST) offers an optimal trade-off between denoising and detail preservation.
+To further adapt WST to sonar-specific patterns, we introduce learnable scale and orientation offsets $(\Delta j,\Delta\theta)$, enabling dynamic deformation of the original Morlet wavelets:
+$\psi_{j,\theta}(x)\;\longmapsto\;\psi_{j+\Delta j,\,\theta+\Delta\theta}(x),\quad x\in\mathbb{R}^{2},$ (2)
+where $j\in\{1,\dots,J\}$ and $\theta\in\{\theta_{1},\dots,\theta_{K}\}$, and each $\psi$ is centered on pixel coordinate $x$.
+The first-order deformable scattering coefficients are defined by
+$S_{1}(I;j,\theta)=\bigl{\lvert}\,I\ast\psi_{\,j+\Delta j,\,\theta+\Delta\theta}\bigr{\rvert}\ast\phi.$ (3)
+[…]
+Second-order deformable scattering coefficients further capture cross-scale interactions (computed only for $j_{1}<j_{2}$ or $\theta_{1}<\theta_{2}$ to avoid duplication):
+$S_{2}(I;$ $\,j_{1},\theta_{1},j_{2},\theta_{2})$ (4)
+$=\bigl{|}\,|I\ast\psi_{j_{1}+\Delta j_{1},\theta_{1}+\Delta\theta_{1}}|$
+$\quad\;\ast\psi_{j_{2}+\Delta j_{2},\theta_{2}+\Delta\theta_{2}}\bigr{|}\ast\phi,$
+Defining $S_{0}=I\ast\phi$, we concatenate all orders into the unified deformable WST feature tensor:
+$G_{\mathrm{WST}}(I)=\mathrm{concat}\bigl{(}S_{0},\,S_{1},\,S_{2}\bigr{)},$ (5)
+where ‘concat’ denotes channel-wise concatenation, yielding an output of size $H\times W\times\bigl{(}1+JK+\binom{JK}{2}\bigr{)}$.
+This tensor is then downsampled by a factor of 2 (bilinear interpolation) and channel-normalized via BatchNorm.
+Our implementation employs $J=3$ scales and $K=6$ orientations, generating $3\!\times\!6=18$ first-order and $\binom{18}{2}=153$ second-order channels.
+Crucially, all offsets $(\Delta j,\Delta\theta)$ are learned jointly with the multi-frame fusion network.
+As illustrated in Fig.
+1(a), the original forward-looking sonar image (red) and multispectral remote sensing image (green) occupy distinct pixel domains, creating a cross-modal degradation gap.
+Conventional approaches attempt to narrow this gap by concatenating handcrafted feature maps with the raw image and applying a fixed $1\times 1$ convolution.
+In contrast, our Deformable WST Feature Bridge (Fig.
+1(b)) directly generates the unified, adaptively optimized feature tensor $G_{\mathrm{WST}}(I)$, shifting data into the yellow feature space.
+This dynamic feature adaptation aligns with the reference-free fusion objective, substantially reducing the degradation gap and improving the transferability of pretrained remote-sensing models for sonar image enhancement tasks.
 III-C Multi-Frame Image Fusion Network
 Figure 3: Illustration of the multi-frame fusion network.
-The proposed network recursively fuses aligned forward-looking sonar frames in a pairwise manner—each step combines the current state with a new frame—until all K frames are aggregated.
-A median reference image guides self-supervised enhancement through downsampling and gradient consistency losses, enabling effective speckle noise suppression and high-resolution output (from 376×376 to 800×800).
-Forward-looking sonar systems typically capture images at high frame rates, which allows for the effective use of complementary information between consecutive frames to naturally eliminate speckle noise.
-As shown in Figure 4, our designed multi-frame image fusion network does not require an additional registration module because the collected sonar images come with accurate pose information.
-The overall architecture comprises three modules: encoding, fusion, and decoding. Let the input low-resolution image sequence be denoted by
-$J=\{J(1),J(2),\ldots,J(K)\}$ (4)
-where ${J^{(i)}}$ represents the $i^{\text{th}}$ frame.
-The encoding module ${E(\cdot)}$ extracts features from each frame, mapping the sequence to a high-dimensional feature space; the fusion module ${F(\cdot)}$ integrates these features into a global encoding ${Z=F(E(J))}$ and the decoding module ${D(\cdot)}$ upsamples and reconstructs the global encoding ${Z}$ to generate the final super-resolution image ${Y}$, namely
-$Y=D(F(E(J)))$ (5)
-Since it is challenging to acquire high-quality sonar images in real underwater scenarios, we follow [25] and adopt a downsampling mean squared error (MSE) loss (denoted as ${{\cal L}_{DS}}$) for self-supervised training.
-This loss ensures that the enhanced super-resolution image remains globally consistent, on a pixel level, with the original low-resolution image.
-Let the downsampling function be ${H(\cdot)}$, and define ${J^{*}}$ as the median image, which must satisfy:
-$H(Y)\approx J^{*}$ (6)
-$J^{*}(u,v)=\operatorname{median}\{J^{(1)}(u,v),J^{(2)}(u,v),\ldots,J^{(M)}(u,v)\}$ (7)
-where ${(u,v)}$ denotes the pixel coordinates in the image. Optimization is performed using the downsampling MSE loss:
-$L_{down}=\|H(Y)-J^{*}\|_{2}^{2}$ (8)
-However, because the downsampling loss primarily focuses on global pixel information, it tends to overly smooth high-frequency details, thereby weakening key features in target regions and affecting subsequent target detection.
-To address this issue, we introduce a local consistency loss and a gradient consistency loss to ensure that the enhanced super-resolution image preserves the local variation trends of the low-resolution image.
-By comparing the differences in adjacent pixel regions, we can effectively retain high-frequency details and avoid over-smoothing:
-$L_{\mathrm{con}}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\left(\left|Y_{p}-Y_{q}\right|-\left|J_{p}^{*}-J_{q}^{*}\right|\right)$ (9)
-$L_{\mathrm{grad}}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\left(\left|\nabla(Y_{p})-\nabla(Y_{q})\right|-\left|\nabla(J_{p}^{*})-\nabla(J_{q}^{*})\right|\right)$ (10)
-[…]
-The final loss function ${L_{\mathrm{total}}}$ is expressed as:
-$L_{\rm total}=L_{\rm down}+\lambda_{\rm con}L_{\rm con}+\lambda_{\rm grad}L_{\rm grad}$ (11)
-where ${\lambda_{\mathrm{con}}}$ and ${\lambda_{\mathrm{grad}}}$ are the hyperparameters that balance each component of the loss function.
+The network recursively fuses aligned forward-looking sonar feature tensors in a pairwise manner until all $K$ frames are aggregated.
+A median reference image guides reference-free enhancement through downsampling and gradient consistency losses, enabling effective speckle noise suppression and high-resolution output (from $376\times 376$ to $800\times 800$).
+Forward-looking sonar systems typically capture images at high frame rates, enabling complementary information across consecutive frames to naturally suppress speckle noise.
+In our pipeline, each raw frame $I^{(i)}\in\mathbb{R}^{H\times W}$ is first processed by the Deformable WST Feature Bridge to produce a unified feature tensor
+$\tilde{J}^{(i)}=G_{\mathrm{WST}}\bigl{(}I^{(i)}\bigr{)},\quad i=1,\dots,K.$
+We then form the input sequence
+$J=\{\tilde{J}^{(1)},\tilde{J}^{(2)},\ldots,\tilde{J}^{(K)}\}.$
+The network comprises three modules: encoding $E(\cdot)$, fusion $F(\cdot)$, and decoding $D(\cdot)$:
+$Z$ $=F\bigl{(}E(J)\bigr{)},$ (6)
+$Y$ $=D(Z),$ (7)
+where $Y\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ is the super-resolved output.
+For reference-free training, we adopt a downsampling mean squared error loss. Let $H(\cdot)$ denote bicubic downsampling by factor 2, and define the pixel-wise median reference
+$J^{*}(u,v)=\operatorname{median}\bigl{\{}J^{(i)}(u,v)\ \big{|}\ i=1,\dots,K\bigr{\}}.$
+The downsampling loss is
+$L_{\rm down}\;=\;\bigl{\lVert}H(Y)-J^{*}\bigr{\rVert}_{2}^{2}.$ (8)
+To preserve high-frequency details, we add a local consistency loss
+$L_{\rm con}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\Bigl{(}\lvert Y_{p}-Y_{q}\rvert-\lvert J^{*}_{p}-J^{*}_{q}\rvert\Bigr{)}$ (9)
+and a gradient consistency loss
+$L_{\rm grad}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\Bigl{(}\lvert\nabla Y_{p}-\nabla Y_{q}\rvert-\lvert\nabla J^{*}_{p}-\nabla J^{*}_{q}\rvert\Bigr{)},$ (10)
+where $p$ indexes local patches, $\sigma(p)$ its neighbors, and $\nabla$ the spatial gradient.
+The overall objective is
+$L_{\rm total}=L_{\rm down}\;+\;\lambda_{\rm con}L_{\rm con}\;+\;\lambda_{\rm grad}L_{\rm grad},$ (11)
+with hyperparameters $\lambda_{\rm con},\lambda_{\rm grad}$ balancing each term.
 IV Experiments
-In this section, we demonstrate that our method outperforms existing approaches [15, 25] on a real-world forward-looking sonar dataset containing three target types.
-Our experiments, conducted on a 24GB NVIDIA GeForce GTX 4090, show that the accurate pose information in the sonar images eliminates the need for additional registration before multi-frame fusion, allowing 16 images to be fused in just 0.3 seconds during inference.
+Most existing sonar image enhancement methods rely on supervised learning with simulated data, which often fails to generalize to real-world scenarios.
+In contrast, we compare against two recent state-of-the-art reference-free techniques—SEGSID [24] and Highres-based method  [25]—both of which do not require high-quality reference images.
+We evaluate all methods on a self-collected forward-looking sonar dataset featuring three target types.
+Our experiments, conducted on a 24 GB NVIDIA GeForce GTX 4090, leverage accurate pose information to avoid any additional registration before multi-frame fusion.
+By closing the gap between sonar and remote-sensing imagery, our Deformable WST Feature Bridge enables the direct fine-tuning of hyperspectral remote-sensing pretrained weights from [46], accelerating convergence and improving enhancement quality on real sonar data.
+Coupled with accurate pose metadata, this adaptation allows the fusion of 16 frames in just 0.3 s at inference.
 IV-A Forward-Looking Sonar Datasets
@@
 Combining this rotation data with the initial sonar pose enables calculation of the sensor’s pose at each instance, eliminating the need for additional image registration prior to multi-frame fusion.
-The dataset comprises multi-frame forward-looking sonar images capturing three target types from various materials and viewpoints: rubber tires (exhibiting severe acoustic absorption), metal torpedo models, and glass-reinforced plastic (GRP) conical frustum.
-Each target class contains 3,200 images grouped into sets of 16 frames.
-We randomly selected 180 groups for training and 10 groups each for validation and testing.
-All images have been converted from polar to Cartesian coordinates at a resolution of 376 × 376.
+Each target type comprises 3,200 sonar images grouped into sequences of 16 frames each.
+We randomly selected 180 groups for training, and reserved 10 groups each for validation and testing.
+All images were converted from polar to Cartesian coordinates at a resolution of 376 × 376 pixels.
+Moreover, instead of training our fusion network entirely from scratch on sonar data, we initialize it with the pretrained weights provided by [25], thereby leveraging existing remote‐sensing models for faster convergence and improved generalization.
+Despite being acquired in a controlled environment, our sonar images exhibit severe degradation reflective of realistic underwater scenarios, including multipath scattering, acoustic absorption, and intrinsic speckle noise.
+The dataset features three targets representing typical degradation conditions: rubber tires with severe acoustic absorption causing low contrast and blurred edges; metal torpedo models exhibiting complex multipath reflections and intense speckle noise; and a glass-reinforced plastic (GRP) conical frustum with moderate acoustic scattering characteristics commonly encountered in underwater missions.
+A circular trajectory ensuring full 360° coverage around each target was adopted, providing diverse angular views and accurate pose estimation without additional image registration, essential for evaluating multi-frame fusion robustness.
+Thus, our dataset effectively represents real-world sonar imaging challenges, validating the robustness and applicability of our enhancement method.
 IV-B Evaluation Metrics for Enhancement Performance
-In underwater environments, high-quality reference images are scarce, making common metrics for optical image enhancement, such as PSNR, unsuitable.
-We adopt methods from [24, 25] to evaluate the denoising performance and the preservation of high-frequency details in sonar image enhancement.
-Standard deviation (STD) is used to assess image smoothness—lower STD indicates stronger noise suppression—while average gradient (AG) measures high-frequency detail, with higher AG reflecting richer detail.
+In underwater environments, high‐quality reference images are scarce, making common metrics for optical image enhancement (e.g., PSNR, SSIM) impractical.
+Instead, we follow established reference-free sonar enhancement works [24, 25] and employ two no‐reference metrics that capture complementary aspects of image quality:Standard deviation (STD) is used to assess image smoothness—lower STD indicates stronger noise suppression—while average gradient (AG) measures high-frequency detail, with higher AG reflecting richer detail.
 $STD=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(I_{i}-\mu\right)^{2}}$ (12)
@@
 Our method consistently achieves the best results in both standard deviation (STD) and average gradient (AG), demonstrating strong capabilities in suppressing noise and preserving structural details.
-For the torpedo model, the STD is reduced from 17.02 to 13.14, surpassing wavelet-based (21.59) and HighRes-based (14.58) methods.
+For the torpedo model, the STD is reduced from 17.02 to 13.14, surpassing SEGSID (21.59) and HighRes-based (14.58) methods.
 AG improves from 4.93 to 5.92, indicating sharper edges and clearer target contours.
@@
 (a) Torpedo model
-Fusion Method Original Wavelet [15] Highres [25] Proposed
+Fusion Method Original SEGSID [24] Highres [25] Proposed
 STD $\downarrow$ 17.02 21.59 14.58 13.14
@@
 (b) Tire
-Fusion Method Original Wavelet [15] Highres [25] Proposed
+Fusion Method Original SEGSID [24] Highres [25] Proposed
 STD $\downarrow$ 58.79 45.96 48.80 37.68
@@
 (c) Conical frustum
-Fusion Method Original Wavelet [15] Highres [25] Proposed
+Fusion Method Original SEGSID [24] Highres [25] Proposed
 STD $\downarrow$ 30.41 30.42 28.83 23.24
@@
 FLR + WST 13.143 5.918
+Figure 6: Visualization of handcrafted features. WST is selected for its balanced performance in denoising and detail retention.
 Figure 7: Comparison of enhancement results using Canny and WST transformations.
@@
 IV-D Ablation Study
-To further assess the contribution of the proposed feature space transformation module, we conducted ablation studies within our multi-frame self-supervised enhancement framework, incorporating several widely used feature operators in sonar image processing.
-These include Histogram of Oriented Gradients (HOG), Canny edge detection, Gray-Level Co-occurrence Matrix (GRE), Haar wavelet features (HAAR), and Wavelet Scattering Transform (WST).
+Figure 6 shows the handcrafted feature operators (HOG, Canny, GRE, HAAR, WST).
+We perform ablation experiments within our multi-frame reference-free enhancement framework to demonstrate that WST achieves the best balance between noise suppression and detail preservation, justifying its selection as the primary feature operator.
 Figure 8: Enhancement results across varying sonar viewpoints.
@@
 (c) The proposed method enhances brightness and edge clarity, achieving robust performance under viewpoint changes.
-Table II reports the quantitative results using standard deviation (STD) and average gradient (AG).
-The results highlight notable performance differences.
-Canny achieved the lowest STD, indicating effective noise suppression.
-However, its low AG suggests over-smoothing and the loss of critical high-frequency edge details.
-In contrast, WST delivered the most balanced performance, combining strong noise reduction with superior detail preservation.
-This is largely due to its multi-scale analysis capability, which enables robust extraction and retention of structural features, particularly in low-quality forward-looking sonar images.
-Figure 7 further visualizes the enhancement results of Canny and WST.
-The WST-based output exhibits higher overall intensity, clearer boundaries, and richer detail, while the Canny-based result shows evident blurring and loss of edge information.
-Such high-frequency detail and contrast enhancement are essential for accurate underwater target detection.
-In conclusion, the ablation study demonstrates that WST provides the optimal trade-off between denoising and detail retention, making it the most effective operator for bridging the cross-modal degradation gap.
-As such, WST is selected as the default feature transformation module in our final approach.
+Table II reports quantitative results in terms of standard deviation (STD, lower is better for noise suppression) and average gradient (AG, higher is better for detail preservation).
+While Canny achieves the lowest STD—indicating strong noise removal—its AG is drastically lower than other methods, reflecting significant loss of edge information.
+In contrast, WST attains a near-optimal STD reduction and the highest AG, striking the best balance between denoising and detail retention.
+Figure 7 illustrates this: WST enhances intensity and edge clarity, whereas Canny’s output, despite being smoother, suffers from blurred boundaries.
+Hence, although Canny suppresses noise more aggressively, WST’s balanced performance makes it the preferred choice for our feature transformation module.
 V Conclusion
-This paper presents a self-supervised enhancement method for forward-looking sonar images, combining feature space transformation with multi-frame fusion.
-The proposed transformation module effectively bridges the cross-modal degradation gap between sonar and other domains (e.g., multispectral remote sensing), enabling a practical transfer learning strategy.
-A self-supervised fusion network leverages the high frame rate of sonar systems to integrate complementary frame information, naturally suppressing noise while preserving high-frequency details.
-Experiments on three real-world, heavily degraded sonar datasets show that the method outperforms existing approaches in both quantitative metrics and visual quality.
-With an inference time of 0.3 seconds for 16 frames, the method is also efficient and well-suited for deployment on underwater vehicles to enhance target detection.
+We propose a reference-free framework for forward‐looking sonar enhancement that integrates a Deformable WST Feature Bridge with an end‐to‐end multi‐frame fusion network.
+The Deformable WST bridge uses learnable wavelet scale and orientation offsets to align sonar features with pretrained remote‐sensing distributions, closing the cross‐modal gap.
+The fusion network then aggregates complementary frames to suppress speckle noise and boost target brightness—without any external ground truth.
+On three self‐collected datasets (rubber tire, metal torpedo, GRP frustum), our method outperforms all reference‐free baselines in STD, AG, and visual quality.
+By narrowing the degradation gap, it also allows direct fine‐tuning of remote‐sensing pretrained weights for faster convergence and stronger generalization.
+Achieving 16‐frame processing in just 0.3 s on a single GPU, it is well suited for real‐time underwater deployment.
+Future work will validate the approach in field environments and embed it within autonomous underwater vehicles.
 VI Acknowledgment